OpenStudy (anonymous):

How can exopential equations with unequal bases be solved Ex:64=3^(4x)

OpenStudy (anonymous):

\(64\) has no base

OpenStudy (campbell_st):

well take the log and solve

OpenStudy (anonymous):

\[b^x=A\iff x=\frac{\log(A)}{\log(b)}\] so \[3^{4x}=64\iff 4x=\frac{\log(64)}{\log(3)}\]

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

Is that it or are their more steps?